OpenStudy (anonymous):

y=(x+5)/(3x-15)-5/(x-5)-2/3

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

find all values of x such that y=0

OpenStudy (anonymous):

find the common denominator first.

OpenStudy (anonymous):

would it be 3x^2-30+75

OpenStudy (anonymous):

no

OpenStudy (anonymous):

It is 3(x-5), then follow the usual method

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

Look carefully the denominators of each term, what the common expression their?

OpenStudy (anonymous):

but only the first 2 of them have x in them i see no common expression

OpenStudy (anonymous):

sorry I am not mean to that but to say common factor

OpenStudy (anonymous):

i am very confused

OpenStudy (anonymous):

can I wright the give expression as \[\frac{ x+5 - 15 - 2(x-5) }{ 3(x-5) }\]?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can I?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thus what is 3(x-5)

OpenStudy (anonymous):

the common factor?

OpenStudy (anonymous):

that is what I am trying to explain. Now, what is the simplified form of the above expression?

OpenStudy (anonymous):

-x/3(x-5)

OpenStudy (anonymous):

right -x/3(x-5) = 0 Then use criss cross method

OpenStudy (anonymous):

what is criss cross?

OpenStudy (anonymous):

multiply -x with 1 and 3(x-5) with 0

OpenStudy (anonymous):

why? i will get -x

OpenStudy (anonymous):

because the value of 3(x-5) is does not make the expression -x/3(x-5) zero. Which means the value of -x/3(x-5) to be zero 3(x-5) does not have any roll

OpenStudy (anonymous):

so then y cannt = 3(x-5)

OpenStudy (anonymous):

no to be y = 0, x must be zero

OpenStudy (anonymous):

ok thanks!

OpenStudy (phi):

y=(x+5)/(3x-15)-5/(x-5)-2/3 solve for y=0 \[ \frac{x+5}{3x-15} -\frac{5}{x-5}-\frac{2}{3}= 0\] The first thing you should look for (hope for!) is that the complicated 3x-15 can factor into something simpler. You notice 3 divides evenly into 3x and into 15, so factor out 3: 3(x-5) \[ \frac{x+5}{3(x-5)} -\frac{5}{x-5}-\frac{2}{3}= 0\] now you can find a common denominator, or you can "clear the denominator" Let's clear the bottom: first multiply all terms on both sides of the = sign by 3(x-5). This will "cancel" the bottom in the first term. \[\cancel{ 3(x-5)}\cdot \frac{x+5}{\cancel{3(x-5)}} -3\cancel{(x-5)}\cdot \frac{5}{\cancel{x-5}}-\cancel{3}(x-5)\cdot \frac{2}{\cancel{3}}= 0\]

OpenStudy (phi):

that mess simplifies to \[ x+5 -3 \cdot 5 -(x-5)\cdot 2= 0\] -3*5 is 15, and -(x-5)*2 is (-x+5)*2, or -2x+10: \[ x+5-15 -2x+10 =0 \] finally, 5-15+10= 0, and x-2x is -x: -x=0 or x=0

OpenStudy (phi):

-3*5 is -15, and -(x-5)*2 is (-x+5)*2, or -2x+10:

OpenStudy (anonymous):

so the solution set would be 0,-0 ?

OpenStudy (phi):

normally people do not attach a - to 0. it is just 0. if you are asking what x,y pair is a solution then (0,0) is the answer